OpenStudy (anonymous):

The quantity of 4 times x minus 10 all divided by negative 2

OpenStudy (anonymous):

you mean 4*(-10)/-2 ?

OpenStudy (anonymous):

in that case 20 is the right answer

OpenStudy (theviper):

\[\Large{\frac{4\times x-10}{-2}}\] \[\Large{\frac{4x-10}{-2}}\] \[\Large{\frac{-2(-2x+5)}{-2}=-2x+5}\]

OpenStudy (theviper):

@vacuole u forgot x :\

OpenStudy (anonymous):

i just not used to english pronounce of mathematic things :-p @TheViper

OpenStudy (theviper):

LOL

OpenStudy (anonymous):

you both were great help thank you...and your better then me cause i suck